Algebra

If A3B-AC=AA9 the find the A,B&C if each of them is distinctly different digits.

RUKHMUDDIN RAJGOLI
7 Years agoGrade

1 Answer

Arun
7 Years ago
1. [a3b] - [ac] = [aa9] 

100a + 30 + b - 10a - c = 100a + 10a + 9 
21 + b = 20a + c 

So clearly we need a=1, and that means c = b+1. There are multiple solutions to this problem. 

(138 - 19 = 119, 137 - 18 = 119, etc.)